# MyFramelinecolor=blue, outerlinewidth=0.7pt, roundcorner=12pt, innertopmargin=.5nerbottommargin=.5nerrightmargin=1pt, innerleftmargin=1pt, backgroundcolor=white

### v1

@@ 4.1. Analytification of the intrinsic mirror family
A priori, $\mathcal{X}$ is only a scheme of finite type over $\operatorname{Spec}\mathbb{C}\llbracket t\rrbracket$. But the relations in Proposition 3.3 are indeed polynomial. Thus $\mathcal{X}$ is the base change to $\operatorname{Spec}\mathbb{C}\llbracket t\rrbracket$ of a flat scheme over $\mathbb{A}^{1}$. Denote by
@@
In the rest of this section we work in the analytic category and omit the subscript “$\mathrm{an}$”. Thus from now on, $\mathcal{X}$ is a complex analytic space with two holomorphic functions $\pi,W:\mathcal{X}\to\mathbb{C}$. The additional holomorphic function $w$ is defined on an open subset of $\mathcal{X}$ containing $\mathcal{X}_{0}\setminus X_{\sigma_{0}}$ and is a function in $W$ and $t$. We assume $W$ and $w$ are real for the given real structure on $\mathcal{X}$ and agree with the analytifications of the regular functions denoted by the same symbols on $\mathcal{X}_{I}$ and $\mathcal{X}_{I}\setminus X_{\sigma_{0}}$ for $I=(t^{k+1})$ and some fixed $k\gg 0$. These properties are automatic in our case, and in general can be achieved by an appropriate choice of analytic approximations. We are going to check identities between holomorphic functions up to holomorphic multiples of $t^{k+1}$ and then take the limit $k\to\infty$.
@@ 4.2. Positive real locus
The positivity of order zero slab functions implies that $\mathcal{X}_{0}^{\mathbb{R}}=\mathcal{X}_{0}\cap\mathcal{X}^{\mathbb{R}}$ has a distinguished connected component $\mathcal{X}_{0}^{>}$ that on each irreducible toric component $X_{\sigma_{i}}$ restricts to the positive real locus, the closure of $\mathbb{R}_{>0}^{2}\subseteq(\mathbb{C}^{*})^{2}$ in $X_{\sigma_{i}}$.
@@
Dehomogenizing (4.1) by $U$ shows that a dense set of $\mathcal{X}_{t}$ is isomorphic to $(\mathbb{C}^{*})^{2}$ with coordinates $x,y$, and $W=t\cdot\big{(}x+y+1/(xy)\big{)}$. From this description the statement is immediate by direct computation already done in §2.2. The critical points of $W$ are at $x=y$, $x^{3}=1$, and hence $3t$ is the only real critical value of $W$. ∎
@@
The homeomorphism of $\mathcal{X}^{>}$ with $B$ extends to small $t>0$ away from a neighborhood $U\subset B$ of a codimension two locus in $B$, as one can see by working with the Kato-Nakayama space of $\mathcal{X}_{0}$ as a log space [A, §7.1]. Moreover, away from a neighborhood of a codimension two locus, the full real locus $\mathcal{X}_{t}^{\mathbb{R}}$ for $t>0$ small is an unbranched cover of $B\setminus U$. The codimension two locus is the image under $\mu$ of the zero locus of the slab functions, the center points of the edges of $\sigma_{0}$ in the case of the mirror of $(\mathcal{Y},\mathcal{D})$. Under the assumption of the positivity of the lowest order coefficients of the slab functions, $\mathcal{X}_{0}^{>}$ is disjoint from this zero locus. A local monodromy computation about the codimension two locus then shows that the homeomorphism $\mathcal{X}_{0}^{>}\to B$ extends to a family of homeomorphisms $\mathcal{X}_{t}^{>}\to B$. In particular, $\mathcal{X}_{t}^{>}$ is homeomorphic to $\mathbb{R}^{2}$ for all $t\geq 0$, and $\Gamma_{s,t}$ is topologically a family of disks.
@@ 4.3. Integral over a tropical 1-chain and canonical coordinates
We will need another family of 2-chains $\beta_{s,t}$ for small $t>0$, each homeomorphic to a pair of pants fibering over a graph $\beta^{\text{trop}}_{s}$ in $B$. Its construction is modeled after the construction of tropical 1-cycles in [RS]. The graph $\beta^{\text{trop}}_{s}$ features one trivalent vertex, three bivalent and three univalent vertices. The univalent vertices $v_{1},v_{2},v_{3}$ lie on the circle $\mu\big{(}|w|=s\big{)}$ from (4.6), the trivalent vertex $v_{0}$ is contained in the central cell of $B$, and the univalent vertices lie on the edges of $\overline{\sigma}_{0}$ as sketched. [^5] The figure is simplified by showing the affine structure compatible with $\mu$ at the central fiber of the full intrinsic mirror family. The base change from Remark 3.4 has the effect of moving the singularities of the affine structure at the three vertices to the centers of the edges of $\overline{\sigma}_{0}$, and $\beta_{s}^{\operatorname{trop}}$ may have to be deformed accordingly.
@@
The integral follows from the computation in §3.6 of [RS], and notably Equation (3.17) with both the Ronkin function $\mathcal{R}$ and the gluing data $s_{\sigma,\underline{\rho}}$, $s_{\sigma^{\prime},\underline{\rho}}$ trivial. The integral thus becomes a sum over simple contributions from the vertices of $\beta^{\text{trop}}_{s}$. Up to the global factor ${2\pi i}$, the trivalent vertex contributes $\frac{1}{2}$, each bivalent vertex contributes $\log t$ and each univalent vertex contributes $-\log s$. ∎
@@
The first equality follows from the normalization property (4.3). Restricting to the distinguished fiber $s=1$ exhibits $-t^{3}$ as a canonical coordinate. But for fixed $t>0$, one could equally well view $s^{-3}$ as a canonical coordinate in the codomain of $W$.
@@ 4.4. Main computation: Integration over positive real Lefschetz-thimbles
of degree 00 and associated tangent vector $(1,0)$. Noting that each wall function of an unbounded wall is a Laurent polynomial in $w^{-1}$ with coefficients in $\mathbb{C}[t]/(t^{k+1})$, we can then write
@@
Since all these wall and slab functions have a non-zero constant coefficient, they have no zeros on any interval $[s_{0},\infty)$, $s_{0}>0$, as long as $t>0$ is sufficiently small. For $s_{1}>s_{0}>0$ and such $t$ we now define
$\Gamma_{t}(s_{0},s_{1})=\mathcal{X}^{>0}\cap w^{-1}\big{(}[s_{0},s_{1}]).$
@@
in $B$. Indeed, by our definition of $\mu$, the value of $|w|$ on $\mu^{-1}(a)$ equals $a^{1/2}$. Note also that if there exists a positive real Lefschetz thimble $L_{s,t}$, as in the case of the mirror of $(\mathcal{Y},\mathcal{D})$ from §3.1, then
$\Gamma_{t}(s_{0},s_{1})=\Gamma_{s_{1},t}-\Gamma_{s_{0},t}$
@@
Since the $U_{i}$ cover $\mathcal{X}_{0}\setminus X_{\sigma_{0}}$ the $U_{i}$ cover $\Gamma_{t}(s_{0},s_{1})$ for sufficiently small $t>0$. Moreover, $(w,u_{i})$ or $(w,v_{i})$ from above provide real, oriented coordinate systems on $\Gamma_{t}(s_{0},s_{1})\cap U_{i}$. By (4.13) we can also impose the condition $v_{i}\leq 1$, or equivalently
@@
For $i=m$ the monodromy brings in an additional factor $w^{-e}$ and (4.15) holds with
@@
Now $\Omega$ is closed for $t={\mathrm{const}}$, hence exact along the contractible $2$-chain $\Gamma_{s,t}$. Thus by Stokes’ Theorem,
@@
In higher dimensions, an analogous period computation can be done for the intrinsic mirror of a normal crossing degeneration $(\mathcal{Y},\mathcal{D})$ of a Fano manifold $Y$ with smooth anticanonical divisor $D$. The positive real locus in that case has to be replaced by a real one-parameter family of cycles induced by a tropical $1$-cycle as in [RS] in the asymptotic singular affine manifold $B_{\infty}$ from [CPS, Constr. 5.5]. At least in the case that this asymptotic tropical $1$-cycle is the tropicalization of a family of curves $\mathcal{C}\subset\mathcal{D}$, we expect the period integral computes the generating function of logarithmic Gromov-Witten invariants in $Y$ intersecting $D$ only in a point of $C=\mathcal{C}\cap Y$.
@@ 4.5. Corollary: Takahashi’s log mirror symmetry conjecture
To deduce Takahashi’s enumerative mirror conjecture (2.7) from our periods, we simply observe that in the canonical coordinate $q=-e^{I_{1}(Q)}$, Takahashi’s basis $I_{0}(Q)$, $I_{1}(Q)$, $I_{2}(Q)$ of solutions of the Picard-Fuchs equation is the unique tuple of solutions obeying
@@
The first equality is (4.3). The second equality follows from (4.2) by
$-\exp\big{(}\Pi_{1}(1,t)\big{)}=-\exp\big{(}\pi i+3\log t\big{)}=t^{3}$

### v2

@@ 4.1. Analytification of the intrinsic mirror family
A priori, $\mathcal{X}$ is only a scheme of finite type over $\operatorname{Spec}\mathbb{C}\llbracket t\rrbracket$. But the relations in Proposition 3.3 are indeed polynomial. Note that polynomiality holds whenever $\mathcal{D}$ supports an ample divisor, as discussed at the beginning of §3. Thus $\mathcal{X}$ is the base change to $\operatorname{Spec}\mathbb{C}\llbracket t\rrbracket$ of a flat scheme over $\mathbb{A}^{1}$. Denote by
@@
In the rest of this section we work in the analytic category and omit the subscript $\mathrm{an}$”. Thus from now on, $\mathcal{X}$ is a complex analytic space with two holomorphic functions $\pi,W:\mathcal{X}\to\mathbb{C}$. The additional holomorphic function $w$ is defined on an open subset of $\mathcal{X}$ containing $\mathcal{X}_{0}\setminus X_{\sigma_{0}}$ and is a function in $W$ and $t$. We assume $W$ and $w$ are real for the given real structure on $\mathcal{X}$ and agree with the analytifications of the regular functions denoted by the same symbols on $\mathcal{X}_{I}$ and $\mathcal{X}_{I}\setminus X_{\sigma_{0}}$ for $I=(t^{k+1})$ and some fixed $k\gg 0$. These properties are automatic in our case, and in general can be achieved by an appropriate choice of analytic approximations. We are going to check identities between holomorphic functions up to holomorphic multiples of $t^{k+1}$ and then take the limit $k\to\infty$.
@@ 4.2. Positive real locus
The positivity of order zero slab functions implies that $\mathcal{X}_{0}^{\mathbb{R}}=\mathcal{X}_{0}\cap\mathcal{X}^{\mathbb{R}}$ has a distinguished connected component $\mathcal{X}_{0}^{>}$ that on each irreducible toric component $X_{\sigma_{i}}$ restricts to the positive real locus, the closure of $\mathbb{R}_{>0}^{2}\subseteq(\mathbb{C}^{*})^{2}$ in $X_{\sigma_{i}}$. Denote by $\mathcal{X}_{t}^{>}\subset\mathcal{X}_{t}$ the intersection of the connected component of $\mathcal{X}^{\mathbb{R}}$ containing $\mathcal{X}_{0}^{>}$ with $\mathcal{X}_{t}$.
@@
Dehomogenizing (4.1) by $U$ shows that a dense set of $\mathcal{X}_{t}$ is isomorphic to $(\mathbb{C}^{*})^{2}$ with coordinates $x,y$, and $W=t\cdot\big(x+y+1/(xy)\big)$. From this description the statement is immediate by direct computation already done in §$2.$ The critical points of $W$ are at $x=y$, $x^{3}=1$, and hence $3t$ is the only real critical value of $W$. ∎
@@
The homeomorphism of $\mathcal{X}^{>}$ with $B$ extends to small $t>0$ away from a neighborhood $U\subset B$ of a codimension two locus in $B$, as one can see by working with the Kato-Nakayama space of $\mathcal{X}_{0}$ as a log space [A, §7.1]. Moreover, away from a neighborhood of a codimension two locus, the full real locus $\mathcal{X}_{t}^{\mathbb{R}}$ for $t>0$ small is an unbranched cover of $B\setminus U$. The codimension two locus is the image under $\mu$ of the zero locus of the slab functions, the center points of the edges of $\sigma_{0}$ in the case of the mirror of $(\mathcal{Y},\mathcal{D})$. Under the assumption of the positivity of the lowest order coefficients of the slab functions, $\mathcal{X}_{0}^{>}$ is disjoint from this zero locus. A local computation near this codimension two locus shows that the homeomorphism $\mathcal{X}_{0}^{>}\to B$ extends to a family of homeomorphisms $\mathcal{X}_{t}^{>}\to B$. In fact, the arguments of [A] using the Kato-Nakayama space show that near points of $\mathcal{X}_{0}$ away from the zero locus of the slab function, the lift of the real locus to the Kato-Nakayama space maps to $B$ by a local homeomorphism, and this local homeomorphism extends to small $t$. Alternatively, one could argue by an explicit computation in local holomorphic coordinates compatible with the real structure. In any case, $\mathcal{X}_{t}^{>}$ is homeomorphic to $\mathbb{R}^{2}$ for all $t\geq 0$, and $\Gamma_{s,t}$ is topologically a family of disks.
@@ 4.3. Integral over a tropical 1-chain and canonical coordinates
We will need another family of 2-chains $\beta_{s,t}$ for small $t>0$, each homeomorphic to a pair of pants fibering over a graph $\beta^{\text{trop}}_{s}$ in $B$. Its construction is modeled after the construction of tropical 1-cycles in [RS]. The graph $\beta^{\text{trop}}_{s}$ features one trivalent vertex, three bivalent and three univalent vertices. The univalent vertices $v_{1},v_{2},v_{3}$ lie on the circle $\mu\big(|w|=s\big)$ from (4.6), the trivalent vertex $v_{0}$ is contained in the central cell of $B$, and the bivalent vertices lie on the edges of $\overline{\sigma}_{0}$ as sketched. [^5] The figure is simplified by showing the affine structure compatible with $\mu$ at the central fiber of the full intrinsic mirror family. The base change from Remark 3.4 has the effect of moving the singularities of the affine structure at the three vertices to the centers of the edges of $\overline{\sigma}_{0}$, and $\beta_{s}^{\operatorname{trop}}$ may have to be deformed accordingly.
@@
The integral follows from the computation in §3.6 of [RS], and notably Equation (3.17) with both the Ronkin function $\mathcal{R}$, the gluing data $s_{\sigma\!\!,\;\underline{\rho}}$ and $s_{\sigma^{\prime}\!\!\!\!,\;\underline{\rho}}$ trivial. The integral thus becomes a sum over simple contributions from the vertices of $\beta^{\text{trop}}_{s}$. Up to the global factor $2\pi i$}, the trivalent vertex contributes $\frac{1}{2}$, each bivalent vertex contributes $\log t$ and each univalent vertex contributes $\log s$. ∎
@@
The first equality follows from the normalization property (4.3). Restricting to the distinguished fiber $s=1$ exhibits $t^{3}$ as a canonical coordinate. But for fixed $t>0$, one could equally well view $s^{-3}$ as a canonical coordinate in the codomain of $W$.
@@ 4.4. Main computation: Integration over positive real Lefschetz-thimbles
of degree 0 and associated tangent vector $(1,0)$. Noting that each wall function of an unbounded wall is a Laurent polynomial in $w^{-1}$ with coefficients in $\mathbb{C}[t]/(t^{k+1})$, we can then write
@@
Since all these wall and slab functions have a non-zero constant coefficient, they have no zeros on any interval $s_{0},\infty)$, $s_{0}>0$, as long as $t>0$ is sufficiently small. For $s_{1}>s_{0}>0$ and such $t$ we now define
$\Gamma_{t}(s_{0},s_{1})=\mathcal{X}^{>}\cap w^{-1}\big([s_{0},s_{1}]).$
@@
in $B$. Indeed, by our definition of $\mu$, the value of $|w|$ on $\mu^{-1}(a)$ equals $a^{1/2}$. Note also that if there exists a positive real Lefschetz thimble $\Gamma_{s,t}$, as in the case of the mirror of $(\mathcal{Y},\mathcal{D})$ from §3.1, then
(4.13) $\Gamma_{t}(s_{0},s_{1})=\Gamma_{s_{1},t}-\Gamma_{s_{0},t}$
@@
Since the $U_{i}$ cover $\mathcal{X}_{0}\setminus X_{\sigma_{0}}$, the $U_{i}$ cover $\Gamma_{t}(s_{0},s_{1})$ for sufficiently small $t>0$. Moreover, $(w,u_{i})$ or $(w,v_{i})$ from above provide real, oriented coordinate systems on $\Gamma_{t}(s_{0},s_{1})\cap U_{i}$. By (4.14) we can also impose the condition $v_{i}\leq 1$, or equivalently
@@
For $i=m$ the monodromy brings in an additional factor $w^{-e}$ and (4.16) holds with $v_{m+1}=v_{0}$ and
@@
Now for $s_{0}$ with $0<s_{0}<s$ and $t>0$ sufficiently small as in Proposition 4.3, we can decompose $\Gamma_{s,t}=\Gamma_{s_{0},t}+\Gamma_{t}(s_{0},s)$ as singular chains. Since the integral of $\Omega$ over $\Gamma_{s_{0},t}$ does not vary with $s$, we conclude
@@
In higher dimensions, an analogous period computation can be done for the intrinsic mirror of a normal crossings degeneration $(\mathcal{Y},\mathcal{D})$ of a Fano manifold $Y$ with smooth anticanonical divisor $D$. The positive real locus in that case has to be replaced by a real one-parameter family of cycles induced by a tropical $1$-cycle as in [RS] in the asymptotic singular affine manifold $B_{\infty}$ from [CPS, Constr. 5.5]. At least in the case that this asymptotic tropical $1$-cycle is the tropicalization of a family of curves $\mathcal{C}\subset\mathcal{D}$, we expect the period integral computes the generating function of logarithmic Gromov-Witten invariants in $Y$ intersecting $D$ only in a point of $C=\mathcal{C}\cap Y$.
@@ 4.5. Corollary: Takahashi’s log mirror symmetry conjecture
To deduce Takahashi’s enumerative mirror conjecture (2.7) from our periods, note that dehomogenizing the equations for $\mathcal{X}$ in Proposition 3.3 with respect to $U$ and setting $x=X/U$, $y=Y/U$ leads to $W=t\cdot\big(x+y+\frac{1}{xy}\big)$. Thus $W=1$ agrees with the family $\check{E}_{t}^{\circ}$ in (2.2) considered by Takahashi. Next observe that in the canonical coordinate $q=-e^{I_{1}(Q)}$, Takahashi’s basis $I_{0}(Q)$, $I_{1}(Q)$, $I_{2}(Q)$ of solutions of the Picard-Fuchs equation is the unique tuple of solutions obeying
@@
The first equality is (4.3). The second equality follows from Proposition 4.2 by
$\exp\big(\Pi_{1}(1,t)\big)=\exp\big(\pi i+3\log t\big)=-t^{3}$